OpenStudy (waheguru):

How to find the inverse of y= 3x^2-12x+1

OpenStudy (waheguru):

I am tying to find the inverse of this function

OpenStudy (usukidoll):

:S ohhhh umm errr

OpenStudy (sweetburger):

put y in for each x

OpenStudy (sweetburger):

at least i believe thats how you do it

OpenStudy (waheguru):

I know that but to find the inverse you need to factor it... How do you do that???

jimthompson5910 (jim_thompson5910):

First complete the square for x y= 3x^2-12x+1 y-1= 3x^2-12x y-1 = 3(x^2-4x) y-1 = 3(x^2-4x+4-4) ... see note below y-1 = 3((x^2-4x+4)-4) y-1 = 3((x-2)^2-4) y-1 = 3(x-2)^2-3(4) y-1 = 3(x-2)^2-12 y = 3(x-2)^2-12+1 y = 3(x-2)^2-11 Note: I took half of the x coefficient -4 to get -4/2 = -2. Then I squared that result to get (-2)^2 = 4. You then add and subtract 4 to x^2 - 4x to complete the square to get x^2 - 4x + 4 - 4. The trinomial x^2 - 4x + 4 will factor to (x-2)^2 which is a perfect square. So this is how x^2 - 4x turns into (x-2)^2 - 4 ------------------------- So y= 3x^2-12x+1 is equivalent to y = 3(x-2)^2-11

jimthompson5910 (jim_thompson5910):

Now you swap x and y, then solve for y y = 3(x-2)^2-11 x = 3(y-2)^2-11 x+11 = 3(y-2)^2 3(y-2)^2 = x+11 (y-2)^2 = (x+11)/3 y-2 = +-sqrt( (x+11)/3 ) y = 2+-sqrt( (x+11)/3 ) The sign on the square root will depend on what the initial domain of the original function. So if x > 2, then you'll be focusing on the upper half which will mean you use the "plus" (instead of the "minus) So again, if x > 2, then the inverse is y = 2 + sqrt( (x+11)/3 )

OpenStudy (waheguru):

Thanks alot Sir, your explaation made sense to me and I finally got it